OpenStudy (anonymous):

Use (x+3y^2)^14 to answer the following questions: How many terms are there? Is there one middle term or two middle terms? Explain. Which numerical term(s) is (are) the middle term(s)? Find the middle term(s). Find the last term.

OpenStudy (anonymous):

\[(x+3y ^{2})^{14}\] Is what the equation looks like.

OpenStudy (anonymous):

Yes, I know that. It's the rest I am having trouble with.

OpenStudy (anonymous):

15 terms, One middle term because the number of terms is odd, the middle term would the be the 8th term. Got that much figured.

zepdrix (zepdrix):

\[\Large\rm (x+3y^2)^{14}\] Recall that we start counting from ZERO, so yes, while the 8th term is the middle term, that ends up being the 7th when we count from zero, yah? So maybe we'll call the first term: \(\Large\rm T_0\) Then the middle term is: \(\Large\rm T_7\) We use our binomial theorem and get:\[\Large\rm T_7=\left(\begin{matrix}14 \\ 7\end{matrix}\right)x^{14-7}(3y^2)^7\]

zepdrix (zepdrix):

Mmm lemme make sure I'm doing that correctly...

zepdrix (zepdrix):

\[\Large\rm (x+3y^2)^{14}=\sum_{k=0}^{14} \left(\begin{matrix}14 \\ k\end{matrix}\right)x^{14-k}(3y^2)^k\]

OpenStudy (anonymous):

I have a problem in my notes that has 13 terms and the middle term was the 7th term.

zepdrix (zepdrix):

Did it correspond to k=7 though? Or did it correspond to k=6? Or not sure?

zepdrix (zepdrix):

I'm jus saying that our 8th term should correspond to the value k=7. I shouldn't had called it the "7th term".

OpenStudy (anonymous):

I don't know. She never gave us that formula.

zepdrix (zepdrix):

k=0 gives us our 1st term, k=1 our 2nd term, and so on... k=7 our 8th term.

zepdrix (zepdrix):

No? Ohhh my bad :o

zepdrix (zepdrix):

Well you know what the progression of powers looks like? Or no?

OpenStudy (anonymous):

x goes down and y counts up

zepdrix (zepdrix):

Here is a simple example: \[\rm(x+y)^4=\text{_}x^4y^0+\text{_}x^3y^1+\text{_}x^2y^2+\text{_}x^1y^3+\text{_}x^0y^4\]Yes good good.

ganeshie8 (ganeshie8):

yeah k=7 gives the 8th term in expansion

zepdrix (zepdrix):

So notice in this example, the MIDDLE term is the 3rd, but corresponds to a power of 2 on the y, yes?

zepdrix (zepdrix):

In the example I left spaces for the coefficients, I hope that wasn't confusing. Those come from Pascal's Triangle. They're not filled in yet.

OpenStudy (anonymous):

Ok, so the xy in this problem would be \[x ^{7}y ^{7}\]?

zepdrix (zepdrix):

The powers should add up to 14, so yes that looks correct. Our x was x. But in your problem your "y", the thing holding the second slot, was actually 3y^2, yes?

zepdrix (zepdrix):

\[\Large\rm x^7(3y^2)^7\]

OpenStudy (anonymous):

So...\[x ^{7}2187y ^{y}\]?

OpenStudy (anonymous):

Or no, y^14

zepdrix (zepdrix):

\[\Large\rm \text{__}x^7(3y^2)^7\]Ok good :) But let's also leave a slot for the coefficient in front.

zepdrix (zepdrix):

We can either go to Pascal's Triangle and scroll allllll the way down to the 13th row, or we can use the uhhhh.. yah that thing. with the.. factorials, yah?

zepdrix (zepdrix):

\[\Large\rm \text{__}\cdot2187x^7y^{14}\]

OpenStudy (anonymous):

Factorial or Combination?

zepdrix (zepdrix):

Ya :) Combination

OpenStudy (anonymous):

\[\left(\begin{matrix}14 \\ 7\end{matrix}\right)=3432\]

zepdrix (zepdrix):

\[\Large\rm =3432\cdot2187x^7y^{14}\]Ah! Good good good.

OpenStudy (anonymous):

Wait, multiply the 3432 and 2187? wouldn't they just be placed on their respective variables?

zepdrix (zepdrix):

Yes they could stay there :)\[\Large\rm =3432x^7(3y^2)^7\]\[\Large\rm =3432x^72187y^{14}\]We have 4 things multiplying together. It's not just a coefficient attached to the x and another attached to the y. But remember, multiplication is `commutative`, meaning, we can multiply things in any order. So we can choose to bring both coefficients to the front and combine them.

OpenStudy (anonymous):

So I have 7505784x^7y^14?

zepdrix (zepdrix):

yayyy good job \c:/

OpenStudy (anonymous):

Now we get to find the last term. LOL

zepdrix (zepdrix):

So what will the powers be on x and "y"?

zepdrix (zepdrix):

On the last term, your x power will have counted all the way down, while the y power will have counted all the way up, yah?

zepdrix (zepdrix):

Remember that the coefficient in front is always a 1 for the first and last terms.

OpenStudy (anonymous):

No x. Y would be 14

zepdrix (zepdrix):

Ahh I gotta go D: my game starting!

zepdrix (zepdrix):

Good good, power on x is 0, so x^0 = 1. No x appears.

OpenStudy (anonymous):

Nooooo! You're killing me!!!

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

See, I don't know where you are but it is 3 am here.

OpenStudy (anonymous):

and my y was wrong. It would be (y^2)^14 which would be y^28

zepdrix (zepdrix):

"y" should be 14, yes. Maybe I should stop calling it that :d

zepdrix (zepdrix):

Our second term is 14 power, so \(\Large\rm (3y^2)^{14}\)

OpenStudy (anonymous):

I have figured the last term to be 4782969y^28

zepdrix (zepdrix):

Boy that's a big number :o Yah looks correct!

OpenStudy (anonymous):

I know. It has taken me a whole page to work it out.

zepdrix (zepdrix):

WHUT +_+ I hope you're using a calculator! lol

OpenStudy (anonymous):

I had trouble keeping up when taking notes in class so that is why I didn't fully understand it but now I do.

OpenStudy (anonymous):

Almost done. Will show you. And yes, calculator.

OpenStudy (anonymous):

Barely got it all in the frame. Thought it was going to take two pictures.

zepdrix (zepdrix):

You expanded the whole thing? XD Buhahaha! That's too funny.

zepdrix (zepdrix):

Wait wait I have another problem for you to do then:\[\Large\rm (2x+3y)^{73}\]

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

Yeah....no

OpenStudy (anonymous):

And I technically didn't have to expand it but It really helped me to understand how to do it. Now I can finish the assignment.

zepdrix (zepdrix):

Ah cool stuff c:

OpenStudy (anonymous):

Yeah, I guess. LOL Now don't you want to go help me verify my post on the Discrete board?

OpenStudy (bradely):

Step by step answer http://www.mathskey.com/question2answer/23988/trig-question